40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5296     Entry time: Wed Aug 24 11:40:21 2011
Author: jamie, jenne, kiwamu, suresh, steve 
Type: Update 
Category: SUS 
Subject: problem with ITMX 

ITMX was drag wiped, and the suspension was put back into place.  However, after removing all of the earthquake stops we found that the suspension was hanging in a very strange way.

The optic appears to heavily pitched forward in the suspension.  All of the rear face magnets are high in their OSEMs, while the SIDE OSEM appears fine.  When first inspected, some of the magnets appeared to be stuck to their top OSEM plates, which was definitely causing it to pitch forward severely.  After gently touching the top of the optic I could get the magnets to sit in a more reasonable position in the OSEMs.  However, they still seem to be sitting a little high.  All of the PDMon values are also too low:

  nominal now
UL 1.045 0.767
UR 0.855 0.718
LR

0.745

0.420

LL

0.780

0.415
SD

0.840

0.752

Taking a free swing measurement now.

ELOG V3.1.3-